Exercise 5S-7 Equivalent Units of Production and Cost per Equivalent Unit-FIFO Method [LO5S-6, LO5S-7]

Pureform, Inc., uses the FIFO method in its process costing system. It manufactures a product that passes through two departments. Data for a recent month for the first department follow:

Units Materials Labor Overhead
Work in process inventory, beginning 97,000 $ 101,200 $ 54,300 $ 69,600
Units started in process 918,000
Units transferred out 940,000
Work in process inventory, ending 75,000
Cost added during the month $ 1,674,620 $ 850,465 $ 995,225

The beginning work in process inventory was 90% complete with respect to materials and 75% complete with respect to labor and overhead. The ending work in process inventory was 70% complete with respect to materials and 50% complete with respect to labor and overhead.

Required:

1. Compute the first department's equivalent units of production for materials, labor, and overhead for the month.

2. Compute the first department's cost per equivalent unit for materials, labor, overhead, and in total for the month. (Round your answers to 2 decimal places.)
Answer #1

Equivalent unit

Whole unit EUP-Material EUP-Labor EUP-Overhead
Beginning work in process 97000 97000*10% = 9700 97000*25% = 24250 24250
Unit started and completed 940000-97000 843000 843000 843000
Ending WIP 75000 75000*70% = 52500 37500 37500
Equivalent units 905200 904750 904750

2) Cost per equivalent unit

Material Labor Overhead Total
Cost added 1674620 850465 995225 3520310
Equivalent unit 905200 904750 904750
Cost per equivalent unit 1.85 0.94 1.1 3.89